THERMAL DECOMPOSITION OF ACETALDEHYDE STUDIED BY MATRIX IR AND PIMS SPEC-TROSCOPY

\underline{ANGAYLE VASILIOU}, KRZYSZTOF PIECH, G. BARNEY ELLISON, University of Colorado, Boulder, CO 80309-0215; MARK R. NIMLOS, Center for Renewable Chemical, Technologies & Materials, NREL, 1617 Cole Blvd., Golden, CO 80401; JOHN W. DAILY, Department of Mechanical Engineering, University of Colorado, Boulder, CO 80309-0427; JOHN F. STANTON, Institute for Theoretical Chemistry, Department of Chemistry, University of Texas, Austin, TX 78712-0165.

In all previous experimental studies of the thermal decomposition of acetaldehyde (CH_{3}CHO), the products were presumed to be CH_{3} + CHO. These species result from cracking of the weakest bond. Other routes are possible: (DH_{298}(CH_{3}−CHO) = 84.8 ± 0.2 kcal mol^{−1}; DH_{298}(CH_{3}CO-H) = 89.4 ± 0.3 kcal mol^{−1}; DH_{298}(H-CH_{2}CHO) = 92 ± 2 kcal mol^{−1}. This work explores the possibility of other thermal decomposition pathways, that result via C-H bond scission. We have used a resistively heated SiC tubular reactor with a 65 μsec residence time to study the thermal cracking of acetaldehyde. The decomposition products are identified by two independent techniques: 118.2 nm (10.487 eV) VUV photoionization mass spectroscopy and infrared absorption spectroscopy in a cryogenic matrix. The observed dissociation channels seem to be:

CH_{3}CHO + Δ → CH_{3}CO + H
→ CH_{2}=CHOH
→ CH_{2}=C=O
→ CH_{3} + HCO